### 02 Jan 2024 v1

# Quality and Quantity of Machine Translation References for Automated Metrics

Abstract
Automatic machine translation metrics often use human translations to determine the quality system translations. Common wisdom in the field dictates that the human references should be of very high quality. However, there are no cost-benefit analyses that could be used to guide practitioners who plan to collect references for machine translation evaluation.
We find that higher-quality references lead to better metric correlations with humans at the segment-level. Having up to 7 references per segment and taking their average helps all metrics. Interestingly, the references from vendors of different qualities can be mixed together and improve metric success. Higher quality references, however, cost more to create and we frame this as an optimization problem: given a specific budget, what references should be collected to maximize metric success. These findings can be used by evaluators of shared tasks when references need to be created under a certain budget.
1 Introduction
††footnotetext: github.com/ufal/optimal-reference-translationsMachine translation systems are robustly evaluated through human annotation. This is non-scaleable and non-replicable (Freitag et al., 2021a) for settings such as shared tasks where number of teams submit automatic translations of the same testset. Automatic metrics aim to provide a cheap and replicable solution. Given the translation and possibly the source and reference segments, they produce a score that correlates with what a human annotator would predict. There is support and evidence for not using references (Lommel, 2016) in metrics, i.e. quality estimation (Specia and Shah, 2018; Rei et al., 2021). Still, most of the commonly used metrics (Section 2) require human reference translations (Freitag et al., 2023). These metrics work by comparing either the overlap on the surface-level (e.g. BLEU, Papineni et al., 2002) or semantic representations (e.g. Comet, Rei et al., 2020).
Humans also do not always arrive at the perfect translation and thus the quality of the references themselves varies (Castilho et al., 2018). Given some cases of very poor translations, such as non-translation,111Keeping the text untouched in the source language. the reference-based metrics would clearly fail. While low-quality references are known to decrease the metric correlations (Freitag et al., 2023), the extent of this effect and interactions with other phenomena remains unclear. Many automatic machine translation metrics support multiple references for a single translation natively or by using an average. The usefulness of multiple references is disputed (Freitag et al., 2020) for state-of-the-art system evaluation. Additionally, a professional experienced translator is likely to produce a better translation than an average crowd-worker. However, the cost of a high-quality human translation is likely also much higher. In this paper, we aim to quantify the trade-off between reference quality, quantity and cost for segment-level automated metric performance. We base our experiments on a small-scale EnglishCzech dataset with multiple references of varying qualities.
We pose a series of research questions with immediate implications for practitioners who evaluate machine translation systems. The short answers here are only summaries and more elaborate discussions are in their respective sections.
Q:
Are higher-quality references useful? (Sec. 3.1)
A:
Yes. Low-quality references should be avoided. Too much investment has diminishing returns.
Q:
Are multiple references useful? (Sec. 3.2)
A:
Yes. Averaging or taking the maximum across reference improves the metrics.
Q:
How to allocate a reference budget? (Sec. 3.4)
A:
By not focusing exclusively on either quality or quantity of references but their combination. This can be computed by Algorithm 1, given a list of vendors and their attributes.
2 Setup
To evaluate the effect of references on automatic machine translation evaluation, we need data with controlled references and reference-based metrics.
Optimal Reference Translations.
Zouhar et al. (2023) re-evaluated in detail a subset of the EnglishCzech testset from the News domain used in the WMT2020 campaign (Barrault et al., 2020). New data was created by translating the original source in 4 different human settings ranging from generic translation vendors to academic translatologists. This phase was followed up with a human annotation and post-editing phase.
Zouhar et al. (2023) are concerned with whether the human quality of the references is really the highest achievable one. They stop short of evaluating the impact of this on machine translation evaluation. We re-purpose their data and system submissions from Barrault et al. (2020). We refer to the references, from lowest to highest quality of the source, as R1, R2, R3, and R4. See more details in Table 1.
Automated Metrics.
For the metrics, we use BLEU (Papineni et al., 2002), ChrF (Popović, 2015), TER (Snover et al., 2006), Comet (Rei et al., 2020), its referenceless version Comet, and its updated iteration Comet (Rei et al., 2022), and BLEURT (Sellam et al., 2020). We select those as a representative set of widely-used string-matching and trainable metrics. We focus on and evaluate metric success at the segment-level by correlating the metric scores with human scores using Kendall’s .222 This is the standard segment-level evaluation adopted by the WMT metrics shared task (Freitag et al., 2021b, 2022, 2023).
| Source segments & documents | & |
| Average source segment length | tokens |
| Reference segments | |
| Reference post-editing | |
| Systems & system segments | & |
3 Experiments
3.1 Reference Quality is Important
As stated in Section 2, we have access to four human translations of varying quality. In Table 2 we show the metric success measured by correlation with human scores. The metrics stay the same but the references they use are changed. Across both string-matching and parametrized model-based metrics, R1, the worst human translation, leads to the worst metric performance. The best performance is achieved with R3, a standard professional translation. Notably, it is not R4 which was created by professional translatologists and also the most expensive. This can be explained by the presence of translation shifts,333Deviation from the original segment structure or meaning. which occur more frequently on this professionality level, but can negatively impact the utility of the reference (Fomicheva, 2017). See Appendix B for an example and analysis.
| Metric | R1 | R2 | R3 | R4 |
| BLEU | 0.082
0.103
0.109
0.103
| ChrF | 0.090
0.125
0.128
0.123
| TER | 0.082
0.092
0.114
0.105
| COMET | 0.172
0.176
0.185
0.181
| COMET | 0.189
0.195
0.191
0.192
| BLEURT | 0.159
0.156
0.199
0.178
| Average | 0.129
0.141
0.154
0.147
| Comet | 0.171
A simple way of improving a translation is to post-edit (refine) it, which is a cheaper that translating it from scratch (Daems and Macken, 2020). We mark the post-edited version of a reference Rx as Rx and use them as references for the automated metrics in Table 3. The post-editors are either laymen with knowledge of both languages, students of translatology, or professional translators. While the proficiency level plays a role, in most cases the post-edited translations serve as better references.
| Proficiency | R1-R1 | R2-R2 | R3-R3 | R4-R4 |
| Layman | 0.019 | 0.011 | 0.011 | 0.011 |
| Student | 0.009 | 0.005 | 0.001 | 0.002 |
| Professional | 0.025 | 0.011 | 0.004 | 0.002 |
3.2 Multiple References are Useful
The previous section provided an analysis of how individual references affect metric performance. In many situations, however, multiple references are available. While some metrics, such as BLEU, support multiple references natively, one can also aggregate them using either segment-level averages or maximums (i.e. compute multiple scores for each segment and take the average or maximum). In Table 4 we consider three setups: (1) two high-quality references, R3 and R4, (2) all human translations, or (3) all human post-edited translations. Across all metrics, this segment-level aggregation improves the correlation with humans, especially in the case of using the original four human translations. Taking the maximum and not the average has the advantage that there exists a specific reference which yields that particular score.
| Aggregation | R3 | R{3,4} | Rx | Rx |
| Average | 0.154
0.159
0.166
0.164
| Max | 0.154
0.155
0.165
0.167
To systematically study the effect of reference count on metric performance, in Figure 1 we randomly sample references from the whole pool of original and post-edited translations. The biggest gains in metric performance are achieved until seven references and further gains are negligible.
Biggest advantage is gained from at least three references and taking their segment-level average.3.3 References and Metric Scores
To understand the effect of different metrics, we show the average raw scores of each metric in Table 5. While it appears that the higher the raw score, the better the metric performance (low score of R1 and high scores of R3 and R4), this trend does not explain the improvements of using the post-edited versions, e.g. as R1 over R1, or R3 over R3. In fact, the post-edited versions always lead to lower raw scores. This could be the result of further translations shifts as the post-edits are based on a translation and not the source segment.
| Metric | R1 | R2 | R3 | R4 | R1 | R3 |
| BLEU | 24.2 | 31.5 | 37.1 | 27.3 | 23.9 | 31.0 |
| ChrF | 55.7 | 60.3 | 63.0 | 56.1 | 54.5 | 58.4 |
| TER | -63.3 | -53.0 | -48.7 | -59.4 | -64.1 | -58.9 |
| COMET | 65.5 | 68.9 | 68.2 | 61.0 | 60.4 | 61.4 |
| COMET | 84.6 | 84.9 | 84.8 | 83.6 | 83.6 | 83.7 |
| BLEURT | 61.3 | 66.1 | 68.8 | 64.5 | 61.6 | 64.9 |
3.4 Allocating a Budget for References
Usually, it is simple to gather many source sentences and let multiple systems translate them. Evaluating all of them using human annotators is unattainable but running automated metrics is not. However, these require references, which are also costly. It remains unclear how many references and of quality level to ask for given a budget.
Can references be mixed?
To assess what types of configurations of references can lead to an improvement, we first validate if they can be mixed. For example if 25% of the references can come from a cheaper vendor R1 and 75% from a higher-quality but more expensive vendor R2. This is different from the experiment in Table 4 where each segment had exactly two references from the same two sources. In Figure 2 we show certain combinations between the original human translations. Despite the varying absolute scores of metrics under different references (Section 3.3), the mixing of multiple references appears to lead to an almost perfectly linear combination of the endpoint metric performances. This stands in contrast to Mirsarraf et al. (2012), who documented a decade ago the need to normalize the references.
All metrics benefit from higher-quality references, even when they are mixed.Budget Allocation Algorithm.
In Appendix A we approach this task as an optimization problem. We provide a heuristic-based stochastic algorithm to find an assignment of source segments to be translated by vendors of different costs and qualities within a specific budget. For the current dataset, we set the cost of a segment in R1, R2, R3, and R4 to be 1, 1, 2, and 3, respectively. Their quality, a prior knowledge supported by Section 3.1, were set to 1, 2, 4, and 3. The algorithm in Appendix A contains a hyperparameter that controls whether budget will be allocated more towards having multiple references or more towards having fewer but higher-quality references.
In Figure 3 we show ChrF and BLEU correlations when using the references from our algorithm. The optimal preference between quality and quantity changes with increasing budget. Using all of the budget on either increasing quality or quantity would correspond to the bottom or top row, which are not optimal. The best reference configurations for a particular budget, such as , contain a mixture of references from R1, R2, R3, and R4 with multiple references for some segments.
4 Related Work
Reference quality is known to affect machine translation evaluation. Freitag et al. (2023) note that specifically very low-quality references cause problems. This stands in contrast to the pre-neural machine translation era where the reference quality did not play a role (Hamon and Mostefa, 2008). This is likely caused by the much higher quality of systems being compared. Additionally, Freitag et al. (2020) study preventing low-quality references in human translation campaigns.
The BLEU metric (Papineni et al., 2002) was intended to be used with multiple metrics, which is not a common practice anymore. Nevertheless, newer and more sophisticated methods exist to incorporate them (Qin and Specia, 2015). Our results from Figure 1 confirm the older study of Finch et al. (2004) which studies the effect of the reference count on metric performance. Finally, multiple references can be used in training better machine translation systems (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020; Mi et al., 2020) or for analyzing model uncertainty (Ott et al., 2018) or evaluation uncertainty (Zhang and Vogel, 2004, 2010; Fomicheva et al., 2020).
5 Conclusion
We showed that the quality of references is important for accurate automated machine translation metrics. The relationships is not straightforward: translatologists’ translations, despite being the peak human translation quality, are not the best references. Rather, it is the standard commercial professional translation that work best for current metrics. The improvements apply both for string-matching metrics as well as to parametric model-based ones. Taking the average over multiple references provides the biggest benefit, with diminishing returns after 7 references. We also provided a heuristic-based algorithm for finding a good configuration of references given a budget, which surpasses optimizing solely for quantity or quantity.
Future work.
The dataset size prevents system-level investigations. Because there is little point in evaluating segments that are easy to translate, a follow-up approach could prioritize difficult-to translate segments. This is used by Isabelle et al. (2017) for creating a challenge set. Future works should extend this research by asking the question how many segments are needed to fulfill a certain desideratum, such as effect size or metric accuracy.
Limitations
We note the limitation of using only a small dataset and a single language direction due to the costs of creating multiple rounds of high-quality references. However, we are convinced the results hold in other scenarios because the effect directions are the same across multiple metrics and setups.
Input: Source segments , levels , cost function , utility function ,
tradeoff hyperparameter , temperature , budget .
Output: Assignment .
Note: Figure 3 contains a patience mechanism, spans and uses .
Sample a segment to add.
Sample where to move from.
Appendix A Budget Allocation Algorithm
We formalize the problem with a segment cost for a reference on level and the utility . The levels might correspond to translation vendors which have costs and qualities. In Section 3.4, we assumed and . Given a set of source sentences , the goal is to divvy up the segments into different levels. The selection should maximize performance of a particular metric on a number of systems but needs to be under a fixed budget , i.e. . In our setup, to preserve fair comparison, each segment needs to have at least one reference. This is because the smaller the testset, the easier it is to achieve higher but spurious correlations. Therefore, . The formalization explicitly allows for parts of the testset to be translated multiple times.
We present a heuristic algorithm with pseudocode in Algorithm 1 and explanatory illustration of the two operations in Figure 4. The algorithm continually applies one of the two operations until they can either no longer be applied or the budget is reached. The algorithm will always terminate because Add increases the cost and utility and Move increases the utility by at least and , respectively.
In addition to the metric correlations in Figure 3, we show the average number of references per source segment in Figure 5. With focus on quality, each segment has fewer references.
Appendix B Qualitative Analysis
In Table 6 we show a single source segment, system translation and multiple references and the metric scores. BLEU ranges from 0 to 100 and both maximums are almost achieved just with a different reference. The best human translation led to the lowest BLEU score because of a translation shift. This is not surprising because BLEU operates on the surface-level. Unexpectedly, a similar thing happens also with Comet, which uses a distributed semantic representation of the segments. This shows that parametric model-based metrics are not robust to changes in references.
| BLEU | Comet | Reference |
| 10 | 78 | K evropské špičce nově patří i tři skotští studenti |
| 23 | 120 | Tři skotští studenti se umístili mezi nejlepšími v Evropě |
| 23 | 121 | Tři skotští studenti mezi nejlepšími v Evropě |
| 28 | 116 | Tři skotští studenti byli oceněni jako jedni z nejlepších v Evropě |
| 28 | 115 | Tři skotští studenti byli jmenováni jako jedni z nejlepších v Evropě |
| 28 | 114 | Tři skotští studenti byli vyhlášeni jako jedni z nejlepších v Evropě |
| 32 | 117 | Tři skotští studenti byli jmenováni jednimi z nejlepších v Evropě |
| 37 | 122 | Tři skotští studenti byli jmenováni mezi nejlepšími v Evropě |
| 43 | 125 | Tři skotští studenti se zařadili mezi nejlepší v Evropě |
| 43 | 121 | Tři skotští studenti patří mezi nejlepší v Evropě. |
| 43 | 122 | Tři skotští studenti patří mezi nejlepší v Evropě |
| 60 | 127 | Tři skotští studenti zařazeni mezi nejlepší v Evropě |
| 100 | 131 | Tři skotští studenti byli zařazeni mezi nejlepší v Evropě |
Appendix C Implementation Details
BLEU, ChrF and TER used SacreBleu (Post 2018, v:2.3.1) at the segment-level. For the trainable metrics, we use publicly available models.
BLEU: #:1|c:mixed|e:yes|tok:13a|s:exp
ChrF: #:1|c:mixed|e:yes|nc:6|nw:0|s:no
TER: #:1|c:lc|t:tercom|nr:no|pn:yes|as:no
COMET: Unbabel/wmt20-comet-da
COMET: Unbabel/wmt22-comet-da
BLEURT: BLEURT20
| Metric | R1-R1 | R2-R2 | R3-R3 | R4-R4 | |
Layman PE |
BLEU | 0.019 | 0.007 | 0.009 | 0.010 |
| ChrF | 0.027 | 0.015 | 0.016 | 0.019 | |
| TER | 0.026 | 0.015 | 0.013 | 0.014 | |
| COMET | 0.016 | 0.011 | 0.010 | 0.009 | |
| COMET | 0.008 | 0.006 | 0.007 | 0.005 | |
| BLEURT | 0.017 | 0.015 | 0.011 | 0.010 | |
Student PE |
BLEU | 0.010 | 0.001 | 0.004 | 0.001 |
| ChrF | 0.011 | 0.001 | 0.000 | 0.004 | |
| TER | 0.003 | 0.001 | 0.005 | 0.002 | |
| COMET | 0.010 | 0.003 | 0.002 | 0.002 | |
| COMET | 0.002 | 0.001 | 0.000 | 0.002 | |
| BLEURT | 0.021 | 0.022 | 0.005 | 0.004 | |
Professional PE |
BLEU | 0.035 | 0.011 | 0.007 | 0.000 |
| ChrF | 0.040 | 0.010 | 0.008 | 0.004 | |
| TER | 0.023 | 0.014 | 0.003 | 0.002 | |
| COMET | 0.016 | 0.006 | 0.000 | 0.005 | |
| COMET | 0.008 | 0.002 | 0.003 | 0.005 | |
| BLEURT | 0.027 | 0.022 | 0.005 | 0.001 |
| R1 | R2 | R3 | R4 | R{1,2} | R{1,3} | R{1,4} | R{2,3} | R{2,4} | R{3,4} | R{1,2,3} | R{1,2,4} | R{1,3,4} | R{2,3,4} | Rx | ||
Average |
BLEU | 0.082
0.103
0.109
0.103
0.109
0.122
0.114
0.132
0.124
0.114
0.136
0.124
0.125
0.130
0.134
| ChrF | 0.090
0.125
0.128
0.123
0.121
0.135
0.124
0.148
0.139
0.135
0.146
0.135
0.140
0.147
0.147
| TER | 0.082
0.092
0.114
0.105
0.095
0.120
0.107
0.125
0.117
0.120
0.121
0.110
0.123
0.127
0.124
| COMET | 0.172
0.176
0.185
0.181
0.181
0.189
0.185
0.191
0.183
0.188
0.190
0.185
0.190
0.189
0.189
| COMET | 0.189
0.195
0.191
0.192
0.195
0.197
0.194
0.201
0.197
0.195
0.200
0.197
0.197
0.199
0.199
| BLEURT | 0.159
0.156
0.199
0.178
0.171
0.203
0.183
0.201
0.180
0.203
0.201
0.184
0.204
0.202
0.202
| Average | 0.129
0.141
0.154
0.147
0.145
0.161
0.151
0.166
0.157
0.159
0.166
0.156
0.163
0.166
0.166
Max |
BLEU | 0.082
0.103
0.109
0.103
0.116
0.122
0.118
0.135
0.132
0.111
0.138
0.137
0.121
0.135
0.137
| ChrF | 0.090
0.125
0.128
0.123
0.133
0.137
0.129
0.139
0.146
0.135
0.144
0.148
0.140
0.144
0.147
| TER | 0.082
0.092
0.114
0.105
0.101
0.124
0.116
0.132
0.128
0.117
0.132
0.130
0.126
0.135
0.134
| COMET | 0.172
0.176
0.185
0.181
0.177
0.184
0.185
0.180
0.183
0.183
0.181
0.183
0.184
0.181
0.182
| COMET | 0.189
0.195
0.191
0.192
0.195
0.191
0.196
0.191
0.195
0.191
0.192
0.197
0.191
0.192
0.192
| BLEURT | 0.159
0.156
0.199
0.178
0.180
0.199
0.190
0.188
0.181
0.193
0.197
0.192
0.200
0.188
0.197
| Average | 0.129
0.141
0.154
0.147
0.150
0.159
0.156
0.161
0.161
0.155
0.164
0.164
0.160
0.162
0.165
| R1 | R2 | R3 | R4 | Rx | R{1,1} | R{2,2} | R{3,3} | R{4,4} | R{x,x} | ||
Layman PE |
BLEU | 0.101
0.111
0.117
0.113
0.144
0.092
0.107
0.113
0.108
0.140
| ChrF | 0.118
0.139
0.144
0.142
0.164
0.106
0.134
0.137
0.135
0.159
| TER | 0.107
0.107
0.127
0.119
0.142
0.099
0.102
0.123
0.116
0.139
| COMET | 0.188
0.187
0.195
0.190
0.198
0.183
0.184
0.193
0.188
0.197
| COMET | 0.197
0.201
0.198
0.197
0.203
0.195
0.200
0.196
0.196
0.202
| BLEURT | 0.176
0.170
0.210
0.188
0.209
0.169
0.165
0.206
0.186
0.209
| Average | 0.148
0.153
0.165
0.158
0.177
0.141
0.149
0.161
0.155
0.174
Student PE |
BLEU | 0.092
0.104
0.105
0.102
0.123
0.089
0.107
0.108
0.103
0.130
| ChrF | 0.102
0.126
0.128
0.119
0.139
0.097
0.127
0.130
0.122
0.144
| TER | 0.085
0.093
0.119
0.103
0.119
0.084
0.095
0.119
0.104
0.123
| COMET | 0.182
0.179
0.183
0.179
0.186
0.179
0.179
0.186
0.181
0.188
| COMET | 0.191
0.196
0.191
0.189
0.195
0.191
0.197
0.193
0.191
0.197
| BLEURT | 0.180
0.178
0.203
0.174
0.199
0.172
0.170
0.204
0.177
0.202
| Average | 0.139
0.146
0.155
0.144
0.160
0.136
0.146
0.157
0.146
0.164
Professional PE |
BLEU | 0.118
0.114
0.115
0.103
0.127
0.103
0.113
0.113
0.104
0.133
| ChrF | 0.131
0.135
0.136
0.127
0.146
0.113
0.133
0.135
0.126
0.149
| TER | 0.105
0.106
0.116
0.103
0.118
0.094
0.102
0.118
0.104
0.122
| COMET | 0.188
0.182
0.185
0.186
0.190
0.183
0.181
0.189
0.185
0.191
| COMET | 0.198
0.198
0.195
0.196
0.199
0.195
0.199
0.195
0.195
0.200
| BLEURT | 0.186
0.178
0.204
0.176
0.197
0.177
0.172
0.206
0.179
0.202
| Average | 0.154
0.152
0.159
0.149
0.163
0.144
0.150
0.159
0.149
0.166
Acknowledgements
We extend our gratitude to Yasmin Moslem, Isabella Lai, Theia Vogel, and Blaka Sokołowska who proofread this paper in their free time. Ondřej Bojar received funding from Ministry of Education, Youth and Sports of the Czech Republic LM2018101 LINDAT/CLARIAH-CZ and the 19-26934X grant of the Czech Science Foundation (NEUREM3).