### 10 Apr 2024 v5

# Quality and Quantity of Machine Translation References for Automatic Metrics

Abstract
Automatic machine translation metrics typically rely on human translations to determine the quality of system translations. Common wisdom in the field dictates that the human references should be of very high quality. However, there are no cost-benefit analyses that could be used to guide practitioners who plan to collect references for machine translation evaluation. We find that higher-quality references lead to better metric correlations with humans at the segment-level. Having up to 7 references per segment and taking their average (or maximum) helps all metrics. Interestingly, the references from vendors of different qualities can be mixed together and improve metric success. Higher quality references, however, cost more to create and we frame this as an optimization problem: given a specific budget, what references should be collected to maximize metric success. These findings can be used by evaluators of shared tasks when references need to be created under a certain budget.
[…]
Quality and Quantity of Machine Translation
References for Automatic Metrics
Vilém Zouhar Ondřej Bojar |
ETH Zürich Charles University |
vzouhar@ethz.ch bojar@ufal.cuni.cz |
Abstract content
1. Introduction
00footnotetext: github.com/ufal/optimal-reference-translations hf.co/datasets/zouharvi/optimal-reference-translationsMachine translation systems are robustly evaluated through human annotation. This is non-scaleable and non-replicable (Freitag et al., 2021a) for settings such as shared tasks where a number of teams submit automatic translations of the same testset. Automatic metrics aim to provide a cheap and replicable solution. Given the translation and possibly the source and reference segments, they produce a score that correlates with what a human annotator would predict. There is support and evidence for not using references (Lommel, 2016) in metrics, i.e. quality estimation (Specia and Shah, 2018; Rei et al., 2021). Still, most of the commonly used metrics (Section 3) require human reference translations (Freitag et al., 2023). These metrics work by comparing either the overlap on the surface-level (e.g. BLEU, Papineni et al., 2002), of semantic representations (e.g. COMET, Rei et al., 2020) or some downstream task (e.g. MTEQA, Krubiński et al., 2021).
Humans also do not always arrive at perfect translations and thus the quality of the references themselves varies (Castilho et al., 2018). In cases of very poor translations, such as non-translation,111Text left untouched in the source language. the reference-based metrics would clearly fail. While low-quality references are known to decrease the metric correlations (Freitag et al., 2023), the extent of this effect and interactions with other phenomena remains unclear. Many automatic machine translation metrics support multiple references for a single translation natively or by using an aggreation such as the average. For phrase-based MT and BLEU, the trade-off between the number of references vs. the test set size was studied by Bojar et al. (2013, Section 5), concluding that a single-reference test set of 3000 sentences can be comparable to 6–7 references with just 100–200 test sentences. The usefulness of multiple references was later disputed (Freitag et al., 2020) for state-of-the-art system evaluation and some recent metrics do not even support multiple references. Additionally, a professional experienced translator is likely to produce a better translation than an average crowd-worker. However, the cost of a high-quality human translation is likely also much higher.
In this paper, we aim to quantify the trade-off between reference quality, quantity and cost for segment-level automatic metric performance. We base our experiments on a small-scale EnglishCzech dataset with multiple references of varying qualities.
We pose research questions with immediate implications for practitioners. The short answers here are only summaries.
Q:
Are higher-quality references useful for automatic evaluation?
A:
Low-quality should be avoided. Too much investment has diminishing returns. (Sec. 4.1)
Q:
Are multiple references useful?
A:
Yes. Averaging or taking the maximum across reference improves the metrics. (Sec. 4.2)
Q:
How to allocate the budget?
A:
By not focusing exclusively on either quality or quantity of references but their combination. This can be computed by Algorithm 1, given a list of vendors and their attributes. (Sec. 4.4)
2. Related Work
Reference quality is known to affect machine translation evaluation. Freitag et al. (2023) note that very low-quality references reduce metric success. This stands in contrast to the pre-neural machine translation era where the reference quality did not play an important role in certain settings (Hamon and Mostefa, 2008). This is likely caused by the much higher quality of systems being compared. Vernikos et al. (2022) hypothesize that ambiguous and vague references are the culprits of metric success deterioration. Additionally, Freitag et al. (2020) study how to avoid low-quality references in human translation campaigns.
The BLEU metric (Papineni et al., 2002) was intended to be used with multiple metrics, which was only rarely put in practice over the years. Nevertheless, newer and more sophisticated methods exist to incorporate them (Qin and Specia, 2015). Our results from Figure 1 confirm the older observations of Finch et al. (2004) or Bojar et al. (2013, Section 5) who study the effect of the reference count on metric performance. Finally, multiple references can be used in training better machine translation systems (Madnani et al., 2008; Zheng et al., 2018; Khayrallah et al., 2020; Mi et al., 2020) or for analyzing model uncertainty (Ott et al., 2018) or evaluation uncertainty (Zhang and Vogel, 2004, 2010; Fomicheva et al., 2020). It is also used outside of machine translation for measuring consensus (Vedantam et al., 2015).
The budget allocation algorithm is reminiscent of active learning or data selection. In machine translation, this is limited to selecting training examples (Haffari et al., 2009; González-Rubio et al., 2012; van der Wees et al., 2017; Shi and Huang, 2020; Mendonça et al., 2023). We focus on algorithmic data selection for higher-quality evaluation. We aim to complete similar works on practical advice on machine translation evaluation. Kocmi et al. (2021, 2024) study the reliability of metrics from the perspective of deployment decisions. We show that the configuration of references can make these metrics stronger or weaker on segment-level.
3. Setup
To evaluate the effect of references on automatic machine translation evaluation, we need data with controlled references and reference-based metrics.
Optimal Reference Translations.
Zouhar et al. (2023) re-annotate a subset of the EnglishCzech testset from the News domain of the WMT2020 campaign (Barrault et al., 2020). New references were created by translating the original source in 4 different human settings ranging from generic translation vendors to translatology academics following a novel protocol leading to so-called “optimal reference translations” (Kloudová et al., 2023). This phase was followed by a human annotation and post-editing phase performed by 11 annotators of varying professionalities.
Zouhar et al. (2023) study whether the human quality of the references is really the highest achievable one. They stop short of evaluating the impact of this on machine translation evaluation. We re-purpose their data and system submissions from Barrault et al. (2020). We refer to the references, from lowest to highest quality of the source, as R1, R2, R3, and R4. Specifically, R1 to R2 come from standard translation vendors,222Nevertheless, based on observations of Kloudová et al. (2021), R1 are to a large extent post-edits of one of the participating systems. R3 is high-quality translation vendor, and R4 is the work of translatologists (the optimal reference). See Table 1 for basic statistics.
chrF|#:1|c:mixed|e:yes|nc:6|nw:0|s:no
TER|#:1|c:lc|t:tercom|nr:no|pn:yes|as:no
Source segments & documents | & |
Average source segment length | tokens |
Reference segments | |
Reference post-editing | |
Systems & system segments | & |
Automated Metrics.
33footnotemark: 3For the metrics, we use BLEU (Papineni et al., 2002), chrF (Popović, 2015), TER (Snover et al., 2006), COMET (Rei et al., 2020), its referenceless version COMET, and its updated iteration COMET (Rei et al., 2022), and BLEURT (Sellam et al., 2020). We select these as a representative set of widely-used string-matching and trainable metrics.
Metric Evaluation.
We focus on and evaluate metric success at the segment-level (“sentence”-level) by correlating the metric scores with human scores using Kendall’s .444 ; read more on the definition in Macháček and Bojar (2014). Each translation receives a human score and automatic metric scores which are correlated. This is the standard segment-level evaluation adopted by the WMT Metrics Shared Task (Freitag et al., 2021b, 2022, 2023). In our case (WMT2020), the human segment-level judgments were created from Direct Assessment (Graham et al., 2016) judgements following the “daRR” conversion as described by Mathur et al. (2020): Candidate translations from MT systems were scored on their own, independently of other candidates. For each pair of judgements of candidates translating the same source, we construct one golden-truth item of pairwise comparison if the two individual scores differ by more than 25% absolute. As Mathur et al. (2020), we believe that this difference in the judgement is big enough to trust the simulated pairwise comparison.
Metric | R1 | R2 | R3 | R4 |
BLEU | 0.082
0.103
0.109
0.103
chrF | 0.090
0.125
0.128
0.123
TER | 0.082
0.092
0.114
0.105
COMET | 0.172
0.176
0.185
0.181
COMET | 0.189
0.195
0.191
0.192
BLEURT | 0.159
0.156
0.199
0.178
Average | 0.129
0.141
0.154
0.147
COMET | 0.171
![[Uncaptioned image]](x2.png)
Proficiency | R1-R1 | R2-R2 | R3-R3 | R4-R4 |
Layman | 0.019 | 0.011 | 0.011 | 0.011 |
Student | 0.009 | 0.005 | 0.001 | 0.002 |
Professional | 0.025 | 0.011 | 0.004 | 0.002 |
![[Uncaptioned image]](x4.png)
4. Experiments
4.1. Reference Quality is Important
As stated in Section 3, we have access to four human translations of varying quality. In Table 2, we show the metric success measured by correlation with human scores. The metrics stay the same but the references they use are changed. Across both string-matching and parametrized model-based metrics, R1, the worst human translation, leads to the worst metric performance. The best performance is achieved with R3, a standard professional translation. Notably, it is not R4 which was created by professional translatologists and was also the most expensive one. This can be explained by the presence of translation shifts, which occur more frequently on this professionality level, but can negatively impact the utility of the reference (Fomicheva, 2017). Translation shifts in general refer to deviation from the original structure or meaning. For our new references, the translatologists paid attention to preserve the meaning but they often restructured the sentences. They did this to avoid translationese as much as possible and to express the subtleties of information structure (given vs. new information) which is natively expressed via Czech word order. These boosted word order differences make it harder for automatic metrics to match the, rather translationese, candidate and the reference. We anticipate that more fluent large language model-based MT could sound less translationese and the “optimal reference translations” will serve better in this setting. See Section 5 for an example and analysis.
A simple way of improving a translation is to post-edit (refine) it, which is cheaper that translating it from scratch (Daems and Macken, 2020; Zouhar et al., 2021). Moreover, Bojar et al. (2013, Figure 7) show that such post-edited references lead to a better performance of BLEU, because “every n-gram mismatch indicates an error”. With standard references, an n-gram mismatch often means just lack of reference coverage. However, such post-edited references need to be ideally created for each evaluated MT system. In our case, the post-edits were created starting from human reference translations Rx. We mark them Rx and use them as references for the automatic metrics in Table 3. The post-editors are either laymen with knowledge of both languages, students of translatology, or professional translators. While the proficiency level plays a role, in most cases the post-edited translations serve as better references. Table 6 below lists the raw metric score changes in a closer detail.
Aggregation | R3 | R{3,4} | Rx | Rx |
Average | 0.154
0.159
0.166
0.164
Max | 0.154
0.155
0.165
0.167
![[Uncaptioned image]](x6.png)
4.2. Multiple References are Useful
The previous section provided an analysis of how individual references affect metric performance. In many situations, however, multiple references are available. While some metrics, such as BLEU, support multiple references natively, one can also aggregate them using either segment-level averages or maxima (i.e. compute multiple scores for each segment and take the average or maximum). In Table 4 we consider three setups: (1) two high-quality references, R3 and R4, (2) all human translations, Rx, or (3) all post-edited human translations, Rx. Across all metrics, this segment-level aggregation improves the correlation with humans, especially in the case of using the original four human translations. Taking the maximum and not the average has the advantage that there exists a specific reference which yields that particular score. The maximum also reflects the spirit of automated evaluation: measuring some similarity between the candidate and reference translations. With more references, taking the maximum corresponds to first finding the most similar reference. We include all subsets as references in Table 9.
To systematically study the effect of reference count on metric performance, in Figure 1 we randomly sample references from the whole pool of original and post-edited translations, irrespective of their quality. The biggest gains in metric performance are achieved until seven references and further gains are negligible, which is in line with the observations of Bojar et al. (2013, Section 5).
Metric | R1 | R2 | R4 | R3 | R1 | R3 |
BLEU | 24.2 | 31.5 | 27.3 | 37.1 | 23.9 | 31.0 |
chrF | 55.7 | 60.3 | 56.1 | 63.0 | 54.5 | 58.4 |
TER | -63.3 | -53.0 | -59.4 | -48.7 | -64.1 | -58.9 |
COMET | 65.5 | 68.9 | 61.0 | 68.2 | 60.4 | 61.4 |
COMET | 84.6 | 84.9 | 83.6 | 84.8 | 83.6 | 83.7 |
BLEURT | 61.3 | 66.1 | 64.5 | 68.8 | 61.6 | 64.9 |
![[Uncaptioned image]](x11.png)
4.3. References and Metric Scores
To understand the effect of different metrics, we show the average raw scores of each metric in Table 5. While it appears that the higher the raw score, the better the metric performance (low score of R1 and high scores of R3 and R4), this trend does not explain the improvements of using the post-edited versions, e.g. as R1 over R1, or R3 over R3. In fact, the post-edited versions always lead to lower raw scores. This could be the result of either further translation shifts as the post-edits are based on a translation and not the source or additional (fully justified) corrections in the references which lead to fewer matches with the candidates.
4.4. Allocating a Budget for References
Usually, it is simple to gather many source sentences and let multiple systems translate them. Evaluating all of them using human annotators is unattainable but running automatic metrics is not. However, these require references, which are also costly. It remains unclear how many references and of which quality to obtain to achieve the most reliable automatic quality assessment under a given budget.
Input: Source segments , levels , cost function , utility function
, tradeoff hyperparameter , temperature , budget .
Output: Assignment .
Note: Figure 4 contains a patience mechanism instead of exit on error.
Sample where to move from.
Can references be mixed?
To assess what types of configurations of references can lead to the most reliable automatic evaluation, we first validate if references can be meaningfully mixed. For example, if it is viable that 75% of the sources can have references from a cheaper vendor R1 and 25% from a higher-quality but more expensive vendor R3. This is different from Table 4 where each segment had exactly two references from the same two sources. In Section 4.3, we show that using lower-quality references R1 leads to lower absolute metric scores (e.g. BLEU = 24.2) as opposed to higher-quality ones R3 (e.g. BLEU = 37.1). This holds across all metrics. Bojar et al. (2010) observe that lower BLEU scores are less reliable, but they refer to the range of BLEU 20. It is thus questionable if BLEUs at 20–40 correlate differently with human MT quality judgements. In Figure 2, we mix some of the references together for evaluation, but staying at single-reference evaluation. Despite the varying absolute scores of metrics under different references, as explored in Section 4.3, mixing of multiple references leads to an almost perfectly linear combination of the endpoint metric performances. The biggest gains in this respect are obtained by BLEURT, chrF and BLEU, while COMET is almost unaffected. There is no formal guarantee that the mix of score distributions will not lower the overall Kendall’s . Nevertheless, a positive conclusion is that if there is budget to only translate 25% of segments with high quality, it should be done and it can only improve the overall evaluation reliability.
Budget Allocation Algorithm.
We provide a heuristic stochastic algorithm to find an assignment of source segments to be translated by vendors of different costs and qualities within a specific budget. For the current dataset, we set the cost of a segment in R1, R2, R3, and R4 to be 1, 1, 2, and 3, respectively. Their quality (or “fitness” for the purpose of automatic evaluation) were set to 1, 2, 4, and 3 based on our observations in Section 4.1. Algorithm 1 contains a hyperparameter that controls whether the budget will be allocated more towards having multiple references per-segment or more towards having fewer but higher-quality references per-segment, and the temperature than controls the the sampling randomness.
We formalize the problem with a segment cost for a reference on level and the utility . The levels might correspond to translation vendors which have costs and qualities. In our case, and . Given a set of source sentences , the goal is to assign the segments to different levels R1 …R4. The same segment can be assigned to different quality levels at once, leading to multiple references for that segment. The selection should maximize performance of a particular metric on a number of systems but needs to fit under a fixed budget , i.e. . In our setup, to preserve fair comparison, each segment needs to have at least one reference. This is because the smaller the testset, the easier it is to achieve higher but spurious correlations. Therefore, . The formalization explicitly allows for parts of the testset to be translated multiple times but requires the budget to cover at least the full test set with the cheapest references. This requirement can be fulfilled by subsampling the testset, as commonly done in WMT evaluation campaigns (Kocmi et al., 2023, inter alia).
The pseudocode is provided in Algorithm 1 and explanatory illustration of the two operations in Figure 3. The algorithm continually applies one of the two operations until they can either no longer be applied or the budget is reached. The algorithm will always terminate because Add increases the cost and utility and Promote increases the utility by at least and , respectively. Therefore either the budget will be filled or every segment will receive a reference from all vendors.
In Figure 4, we show chrF and COMET correlations when using the references selected by our algorithm. The optimal preference between quality and quantity changes with increasing budget. Using all of the budget on either quality or quantity would correspond to the bottom or top row, which are not optimal. The best reference configurations for a particular budget, such as , four-times the price of the cheapest translation, contain a mixture of references from R1, R2, R3, and R4 with multiple references for some segments. In addition to the metric correlations in Figure 4, we show the average number of references per source segment in Figure 5. With focus on quality, each segment has fewer references.
5. Qualitative Analysis
In Tables 7 and 8, we show a single source segment, one system translation and multiple references and the metric scores. BLEU ranges from 0 to 100 and both extremes are almost achieved just with a different reference. The best human translation led to the lowest BLEU score because of a translation shift. This is not surprising because BLEU operates on the surface-level. Unexpectedly, a similar thing happens also with COMET, which uses a distributed semantic representation of the segments. This shows that parametric model-based metrics are not robust to changes in references. In Table 8, the COMET difference between references is large due to some translators deciding to drop the verb “spolupracovat” (collaborate), which changes the meaning and the system translation is penalized.
6. Conclusion
We showed that the quality of references is important for accurate automatic machine translation metrics. The relationship is not straightforward: translatologists’ translations, despite being the peak translation quality, are not the best references. Rather, it is the standard commercial professional translations that work best for current metrics. The trend applies to both string-matching metrics as well as to parametric model-based ones. Taking the average over multiple references provides the biggest benefit, with diminishing returns after 7 references. We also provided a heuristic-based algorithm for finding a good configuration of references given a budget, which surpasses optimizing solely for quantity or quantity.
Future work.
The dataset size prevents system-level investigations. Because there is little point in evaluating segments that are easy to translate, a follow-up approach could prioritize difficult-to translate segments. This is used by Isabelle et al. (2017) for creating a challenge set. Future works should quantify the references quality and ask how many segments are needed to fulfill a certain desideratum, such as effect size or metric accuracy.
Limitations.
We note the limitation of using a small dataset and a single language translation direction due to the costs of creating multiple rounds of high-quality references. We are convinced the results hold in other scenarios as the effect directions are the same across multiple metrics and setups.
![[Uncaptioned image]](x20.png)
![[Uncaptioned image]](x21.png)
![[Uncaptioned image]](x23.png)
BLEU | COMET | Reference |
10 | 78 | K evropské špičce nově patří i tři skotští studenti |
23 | 120 | Tři skotští studenti se umístili mezi nejlepšími v Evropě |
23 | 121 | Tři skotští studenti mezi nejlepšími v Evropě |
28 | 116 | Tři skotští studenti byli oceněni jako jedni z nejlepších v Evropě |
28 | 115 | Tři skotští studenti byli jmenováni jako jedni z nejlepších v Evropě |
28 | 114 | Tři skotští studenti byli vyhlášeni jako jedni z nejlepších v Evropě |
32 | 117 | Tři skotští studenti byli jmenováni jednimi z nejlepších v Evropě |
37 | 122 | Tři skotští studenti byli jmenováni mezi nejlepšími v Evropě |
43 | 125 | Tři skotští studenti se zařadili mezi nejlepší v Evropě |
43 | 121 | Tři skotští studenti patří mezi nejlepší v Evropě. |
43 | 122 | Tři skotští studenti patří mezi nejlepší v Evropě |
60 | 127 | Tři skotští studenti zařazeni mezi nejlepší v Evropě |
100 | 131 | Tři skotští studenti byli zařazeni mezi nejlepší v Evropě |
![[Uncaptioned image]](x25.png)
BLEU | COMET | Reference |
4 | -42 | Sony a Disney točí třetí film o Spidermanovi |
4 | -33 | Sony a Disney točí třetí film o Spider-Manovi |
8 | -9 | Sony a Disney pracují na třetím filmu o Spider-Manovi |
8 | -4 | Sony a Disney pokračují v práci na třetím filmu o Spider-Manovi |
8 | 1 | Sony a Disney opět pracují na třetím filmu o Spider-Manovi |
8 | 15 | Sony a Disney spolupracují na třetím filmu o Spider-Manovi |
4 | 16 | Sony a Disney budou spolupracovat při natáčení třetího filmo o Spider-manovi |
8 | 28 | Sony a Disney opět spolupracují na třetím filmu o Spider-Manovi |
8 | 30 | Sony a Disney budou spolupracovat na třetím filmu o Spider-Manovi |
8 | 35 | Sony a Disney budou opět spolupracovat na třetím filmu o Spider-Manovi |
17 | 52 | Disney bude znovu spolupracovat se společností Sony na třetím filmu Spider-Man |
10 | 64 | Disney bude se Sony dál pracovat na třetím filmu se Spider-Manem |
50 | 73 | Disney bude spolupracovat se Sony na třetím sólovém filmu o Spider-Manovi |
75 | 99 | Disney se vrací, bude spolupracovat se Sony na třetím filmu o Spider-Manovi |
78 | 106 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém filu Spider-Man. |
79 | 108 | Disney se vrací, bude spolupracovat se Sony na třetím Spider-Man filmu |
100 | 121 | Disney se vrací, bude spolupracovat se Sony na třetím sólovém Spider-Man filmu |
R1 | R2 | R3 | R4 | R{1,2} | R{1,3} | R{1,4} | R{2,3} | R{2,4} | R{3,4} | R{1,2,3} | R{1,2,4} | R{1,3,4} | R{2,3,4} | Rx | ||
Average |
BLEU | 0.082
0.103
0.109
0.103
0.109
0.122
0.114
0.132
0.124
0.114
0.136
0.124
0.125
0.130
0.134
chrF | 0.090
0.125
0.128
0.123
0.121
0.135
0.124
0.148
0.139
0.135
0.146
0.135
0.140
0.147
0.147
TER | 0.082
0.092
0.114
0.105
0.095
0.120
0.107
0.125
0.117
0.120
0.121
0.110
0.123
0.127
0.124
COMET | 0.172
0.176
0.185
0.181
0.181
0.189
0.185
0.191
0.183
0.188
0.190
0.185
0.190
0.189
0.189
COMET | 0.189
0.195
0.191
0.192
0.195
0.197
0.194
0.201
0.197
0.195
0.200
0.197
0.197
0.199
0.199
BLEURT | 0.159
0.156
0.199
0.178
0.171
0.203
0.183
0.201
0.180
0.203
0.201
0.184
0.204
0.202
0.202
Average | 0.129
0.141
0.154
0.147
0.145
0.161
0.151
0.166
0.157
0.159
0.166
0.156
0.163
0.166
0.166
Max |
BLEU | 0.082
0.103
0.109
0.103
0.116
0.122
0.118
0.135
0.132
0.111
0.138
0.137
0.121
0.135
0.137
chrF | 0.090
0.125
0.128
0.123
0.133
0.137
0.129
0.139
0.146
0.135
0.144
0.148
0.140
0.144
0.147
TER | 0.082
0.092
0.114
0.105
0.101
0.124
0.116
0.132
0.128
0.117
0.132
0.130
0.126
0.135
0.134
COMET | 0.172
0.176
0.185
0.181
0.177
0.184
0.185
0.180
0.183
0.183
0.181
0.183
0.184
0.181
0.182
COMET | 0.189
0.195
0.191
0.192
0.195
0.191
0.196
0.191
0.195
0.191
0.192
0.197
0.191
0.192
0.192
BLEURT | 0.159
0.156
0.199
0.178
0.180
0.199
0.190
0.188
0.181
0.193
0.197
0.192
0.200
0.188
0.197
Average | 0.129
0.141
0.154
0.147
0.150
0.159
0.156
0.161
0.161
0.155
0.164
0.164
0.160
0.162
0.165
R1 | R2 | R3 | R4 | Rx | R{1,1} | R{2,2} | R{3,3} | R{4,4} | R{x,x} | ||
Layman PE |
BLEU | 0.101
0.111
0.117
0.113
0.144
0.092
0.107
0.113
0.108
0.140
chrF | 0.118
0.139
0.144
0.142
0.164
0.106
0.134
0.137
0.135
0.159
TER | 0.107
0.107
0.127
0.119
0.142
0.099
0.102
0.123
0.116
0.139
COMET | 0.188
0.187
0.195
0.190
0.198
0.183
0.184
0.193
0.188
0.197
COMET | 0.197
0.201
0.198
0.197
0.203
0.195
0.200
0.196
0.196
0.202
BLEURT | 0.176
0.170
0.210
0.188
0.209
0.169
0.165
0.206
0.186
0.209
Average | 0.148
0.153
0.165
0.158
0.177
0.141
0.149
0.161
0.155
0.174
Student PE |
BLEU | 0.092
0.104
0.105
0.102
0.123
0.089
0.107
0.108
0.103
0.130
chrF | 0.102
0.126
0.128
0.119
0.139
0.097
0.127
0.130
0.122
0.144
TER | 0.085
0.093
0.119
0.103
0.119
0.084
0.095
0.119
0.104
0.123
COMET | 0.182
0.179
0.183
0.179
0.186
0.179
0.179
0.186
0.181
0.188
COMET | 0.191
0.196
0.191
0.189
0.195
0.191
0.197
0.193
0.191
0.197
BLEURT | 0.180
0.178
0.203
0.174
0.199
0.172
0.170
0.204
0.177
0.202
Average | 0.139
0.146
0.155
0.144
0.160
0.136
0.146
0.157
0.146
0.164
Professional PE |
BLEU | 0.118
0.114
0.115
0.103
0.127
0.103
0.113
0.113
0.104
0.133
chrF | 0.131
0.135
0.136
0.127
0.146
0.113
0.133
0.135
0.126
0.149
TER | 0.105
0.106
0.116
0.103
0.118
0.094
0.102
0.118
0.104
0.122
COMET | 0.188
0.182
0.185
0.186
0.190
0.183
0.181
0.189
0.185
0.191
COMET | 0.198
0.198
0.195
0.196
0.199
0.195
0.199
0.195
0.195
0.200
BLEURT | 0.186
0.178
0.204
0.176
0.197
0.177
0.172
0.206
0.179
0.202
Average | 0.154
0.152
0.159
0.149
0.163
0.144
0.150
0.159
0.149
0.166
Acknowledgements
We extend our gratitude to Yasmin Moslem, Isabella Lai, Theia Vogel, Blanka Sokołowska, Raj Dabre, Tom Kocmi, and Farhan Samir, who proofread this paper in their free time. Ondřej Bojar received funding from Ministry of Education, Youth and Sports of the Czech Republic LM2018101 LINDAT/CLARIAH-CZ and the 19-26934X grant of the Czech Science Foundation (NEUREM3).
Bibliographical References
[…]
- Bojar et al. (2010) Ondřej Bojar, Kamil Kos, and David Mareček. 2010. Tackling sparse data issue in machine translation evaluation. In Proceedings of the ACL 2010 Conference Short Papers, 86–91, Uppsala, Sweden. Association for Computational Linguistics.
- Bojar et al. (2013) Ondřej Bojar, Christian Buck, Chris Callison-Burch, Christian Federmann, Barry Haddow, Philipp Koehn, Christof Monz, Matt Post, Radu Soricut, and Lucia Specia. 2013. Findings of the 2013 Workshop on Statistical Machine Translation. In Proceedings of the Eighth Workshop on Statistical Machine Translation, 1–44, Sofia, Bulgaria. Association for Computational Linguistics.
- Castilho et al. (2018) Sheila Castilho, Stephen Doherty, Federico Gaspari, and Joss Moorkens. 2018. Approaches to human and machine translation quality assessment. Translation quality assessment: From principles to practice.
- Daems and Macken (2020) Joke Daems and Lieve Macken. 2020. Post-editing human translations and revising machine translations: Impact on efficiency and quality. In Translation Revision and Post-editing, 50–70. Routledge.
- Finch et al. (2004) Andrew M Finch, Yasuhiro Akiba, and Eiichiro Sumita. 2004. How does automatic machine translation evaluation correlate with human scoring as the number of reference translations increases? In LREC.
- Fomicheva (2017) Marina Fomicheva. 2017. The role of human reference translation in machine translation evaluation. Ph.D. thesis, Universitat Pompeu Fabra.
- Fomicheva et al. (2020) Marina Fomicheva, Lucia Specia, and Francisco Guzmán. 2020. Multi-hypothesis machine translation evaluation. In Proceedings of the 58th Annual Meeting of ACL.
- Freitag et al. (2021a) Markus Freitag, George Foster, David Grangier, Viresh Ratnakar, Qijun Tan, and Wolfgang Macherey. 2021a. Experts, Errors, and Context: A Large-Scale Study of Human Evaluation for Machine Translation. Transactions of the Association for Computational Linguistics, 9:1460–1474.
- Freitag et al. (2020) Markus Freitag, David Grangier, and Isaac Caswell. 2020. BLEU might be guilty but references are not innocent. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 61–71. Association for Computational Linguistics.
- Freitag et al. (2023) Markus Freitag, Nitika Mathur, Chi-kiu Lo, Eleftherios Avramidis, Ricardo Rei, Brian Thompson, Tom Kocmi, Frederic Blain, Daniel Deutsch, Craig Stewart, Chrysoula Zerva, Sheila Castilho, Alon Lavie, and George Foster. 2023. Results of WMT23 metrics shared task: Metrics might be guilty but references are not innocent. In Proceedings of the Eighth Conference on Machine Translation, 578–628. Association for Computational Linguistics.
- Freitag et al. (2022) Markus Freitag, Ricardo Rei, Nitika Mathur, Chi-kiu Lo, Craig Stewart, Eleftherios Avramidis, Tom Kocmi, George Foster, Alon Lavie, and André F. T. Martins. 2022. Results of WMT22 metrics shared task: Stop using BLEU – neural metrics are better and more robust. In Proceedings of the Seventh Conference on Machine Translation, 46–68. Association for Computational Linguistics.
- Freitag et al. (2021b) Markus Freitag, Ricardo Rei, Nitika Mathur, Chi-kiu Lo, Craig Stewart, George Foster, Alon Lavie, and Ondřej Bojar. 2021b. Results of the WMT21 metrics shared task: Evaluating metrics with expert-based human evaluations on TED and news domain. In Proceedings of the Sixth Conference on Machine Translation, 733–774. Association for Computational Linguistics.
- González-Rubio et al. (2012) Jesús González-Rubio, Daniel Ortiz-Martínez, and Francisco Casacuberta. 2012. Active learning for interactive machine translation. In Proceedings of the 13th Conference of the European Chapter of the Association for Computational Linguistics, 245–254. Association for Computational Linguistics.
- Graham et al. (2016) Yvette Graham, Timothy Baldwin, Alistair Moffat, and Justin Zobel. 2016. Can machine translation systems be evaluated by the crowd alone. Natural Language Engineering, FirstView:1–28.
- Haffari et al. (2009) Gholamreza Haffari, Maxim Roy, and Anoop Sarkar. 2009. Active learning for statistical phrase-based machine translation. In Proceedings of Human Language Technologies: The 2009 Annual Conference of the North American Chapter of the Association for Computational Linguistics, 415–423. Association for Computational Linguistics.
- Hamon and Mostefa (2008) Olivier Hamon and Djamel Mostefa. 2008. The impact of reference quality on automatic MT evaluation. In Coling 2008: Companion volume: Posters, 39–42.
- Isabelle et al. (2017) Pierre Isabelle, Colin Cherry, and George Foster. 2017. A challenge set approach to evaluating machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 2486–2496. Association for Computational Linguistics.
- Khayrallah et al. (2020) Huda Khayrallah, Brian Thompson, Matt Post, and Philipp Koehn. 2020. Simulated multiple reference training improves low-resource machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 82–89. Association for Computational Linguistics.
- Kloudová et al. (2023) Věra Kloudová, David Mraček, Ondřej Bojar, and Martin Popel. 2023. Možnosti a meze tvorby tzv. optimálních referenčních překladů: po stopách „překladatelštiny“ v profesionálních překladech zpravodajských textů. Slovo a slovesnost, 84(2):122–156.
- Kloudová et al. (2021) Věra Kloudová, Ondřej Bojar, and Martin Popel. 2021. Detecting post-edited references and their effect on human evaluation. In Proceedings of the Workshop on Human Evaluation of NLP Systems (HumEval), 114–119, Stroudsburg, USA. Association for Computational Linguistics.
- Kocmi et al. (2023) Tom Kocmi, Eleftherios Avramidis, Rachel Bawden, Ondřej Bojar, Anton Dvorkovich, Christian Federmann, Mark Fishel, Markus Freitag, Thamme Gowda, Roman Grundkiewicz, Barry Haddow, Philipp Koehn, Benjamin Marie, Christof Monz, Makoto Morishita, Kenton Murray, Makoto Nagata, Toshiaki Nakazawa, Martin Popel, Maja Popović, and Mariya Shmatova. 2023. Findings of the 2023 conference on machine translation (WMT23): LLMs are here but not quite there yet. In Proceedings of the Eighth Conference on Machine Translation, 1–42, Singapore. Association for Computational Linguistics.
- Kocmi et al. (2021) Tom Kocmi, Christian Federmann, Roman Grundkiewicz, Marcin Junczys-Dowmunt, Hitokazu Matsushita, and Arul Menezes. 2021. To ship or not to ship: An extensive evaluation of automatic metrics for machine translation. In Proceedings of the Sixth Conference on Machine Translation, 478–494. Association for Computational Linguistics.
- Kocmi et al. (2024) Tom Kocmi, Vilém Zouhar, Christian Federmann, and Matt Post. 2024. Navigating the metrics maze: Reconciling score magnitudes and accuracies.
- Krubiński et al. (2021) Mateusz Krubiński, Erfan Ghadery, Marie-Francine Moens, and Pavel Pecina. 2021. Just ask! evaluating machine translation by asking and answering questions. In Proceedings of the Sixth Conference on Machine Translation, 495–506, Online. Association for Computational Linguistics.
- Lommel (2016) Arle Lommel. 2016. Blues for BLEU: Reconsidering the validity of reference-based MT evaluation. Translation Evaluation: From Fragmented Tools and Data Sets to an Integrated Ecosystem, 63.
- Macháček and Bojar (2014) Matouš Macháček and Ondřej Bojar. 2014. Results of the WMT14 metrics shared task. In Proceedings of the Ninth Workshop on Statistical Machine Translation, 293–301, Baltimore, MD, USA. Association for Computational Linguistics.
- Madnani et al. (2008) Nitin Madnani, Philip Resnik, Bonnie J. Dorr, and Richard Schwartz. 2008. Are multiple reference translations necessary? investigating the value of paraphrased reference translations in parameter optimization. In Proceedings of the 8th Conference of the Association for Machine Translation in the Americas, 143–152.
- Mathur et al. (2020) Nitika Mathur, Johnny Tian-Zheng Wei, Markus Freitag, Qingsong Ma, and Ondřej Bojar. 2020. Results of the WMT20 metrics shared task. In Fifth Conference on Machine Translation - Proceedings of the Conference, 688–725, Stroudsburg, PA, USA. Association for Computational Linguistics, Association for Computational Linguistics.
- Mendonça et al. (2023) Vânia Mendonça, Ricardo Rei, Luísa Coheur, and Alberto Sardinha. 2023. Onception: Active Learning with Expert Advice for Real World Machine Translation. Computational Linguistics, 49(2):325–372.
- Mi et al. (2020) Chenggang Mi, Lei Xie, and Yanning Zhang. 2020. Improving adversarial neural machine translation for morphologically rich language. IEEE Transactions on Emerging Topics in Computational Intelligence, (4):417–426.
- Ott et al. (2018) Myle Ott, Michael Auli, David Grangier, and Marc’Aurelio Ranzato. 2018. Analyzing uncertainty in neural machine translation. In Proceedings of the 35th International Conference on Machine Learning, volume 80 of Proceedings of Machine Learning Research, 3956–3965. PMLR.
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of ACL, 311–318. Association for Computational Linguistics.
- Popović (2015) Maja Popović. 2015. chrF: character n-gram F-score for automatic MT evaluation. In Proceedings of the Tenth Workshop on Statistical Machine Translation, 392–395. Association for Computational Linguistics.
- Qin and Specia (2015) Ying Qin and Lucia Specia. 2015. Truly exploring multiple references for machine translation evaluation. In Proceedings of the 18th Annual Conference of the European Association for Machine Translation.
- Rei et al. (2022) Ricardo Rei, José G. C. de Souza, Duarte Alves, Chrysoula Zerva, Ana C Farinha, Taisiya Glushkova, Alon Lavie, Luisa Coheur, and André F. T. Martins. 2022. COMET-22: Unbabel-IST 2022 submission for the metrics shared task. In Proceedings of the Seventh Conference on Machine Translation, 578–585. Association for Computational Linguistics.
- Rei et al. (2021) Ricardo Rei, Ana C Farinha, Chrysoula Zerva, Daan van Stigt, Craig Stewart, Pedro Ramos, Taisiya Glushkova, André F. T. Martins, and Alon Lavie. 2021. Are references really needed? unbabel-IST 2021 submission for the metrics shared task. In Proceedings of the Sixth Conference on Machine Translation, 1030–1040. Association for Computational Linguistics.
- Rei et al. (2020) Ricardo Rei, Craig Stewart, Ana C Farinha, and Alon Lavie. 2020. COMET: A neural framework for MT evaluation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 2685–2702. Association for Computational Linguistics.
- Sellam et al. (2020) Thibault Sellam, Dipanjan Das, and Ankur Parikh. 2020. BLEURT: Learning robust metrics for text generation. In Proceedings of the 58th Annual Meeting of ACL, 7881–7892. Association for Computational Linguistics.
- Shi and Huang (2020) Zhouxing Shi and Minlie Huang. 2020. Robustness to modification with shared words in paraphrase identification. In Findings of the Association for Computational Linguistics: EMNLP 2020, 164–171. Association for Computational Linguistics.
- Snover et al. (2006) Matthew Snover, Bonnie Dorr, Rich Schwartz, Linnea Micciulla, and John Makhoul. 2006. A study of translation edit rate with targeted human annotation. In Proceedings of the 7th Conference of the Association for Machine Translation in the Americas: Technical Papers, 223–231. Association for Machine Translation in the Americas.
- Specia and Shah (2018) Lucia Specia and Kashif Shah. 2018. Machine translation quality estimation: Applications and future perspectives. Translation quality assessment: from principles to practice, 201–235.
- van der Wees et al. (2017) Marlies van der Wees, Arianna Bisazza, and Christof Monz. 2017. Dynamic data selection for neural machine translation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 1400–1410. Association for Computational Linguistics.
- Vedantam et al. (2015) Ramakrishna Vedantam, C Lawrence Zitnick, and Devi Parikh. 2015. Cider: Consensus-based image description evaluation. In Proceedings of the IEEE conference on computer vision and pattern recognition, 4566–4575.
- Vernikos et al. (2022) Giorgos Vernikos, Brian Thompson, Prashant Mathur, and Marcello Federico. 2022. Embarrassingly easy document-level MT metrics: How to convert any pretrained metric into a document-level metric. In Proceedings of the Seventh Conference on Machine Translation (WMT), 118–128. Association for Computational Linguistics.
- Zhang and Vogel (2004) Ying Zhang and Stephan Vogel. 2004. Measuring confidence intervals for the machine translation evaluation metrics. In Proceedings of the 10th Conference on Theoretical and Methodological Issues in Machine Translation of Natural Languages.
- Zhang and Vogel (2010) Ying Zhang and Stephan Vogel. 2010. Significance tests of automatic machine translation evaluation metrics. Machine Translation, 24:51–65.
- Zheng et al. (2018) Renjie Zheng, Mingbo Ma, and Liang Huang. 2018. Multi-reference training with pseudo-references for neural translation and text generation. In Proceedings of the Conference on Empirical Methods in Natural Language Processing, 3188–3197. Association for Computational Linguistics.
- Zouhar et al. (2021) Vilém Zouhar, Martin Popel, Ondřej Bojar, and Aleš Tamchyna. 2021. Neural machine translation quality and post-editing performance. In Proceedings of the 2021 Conference on Empirical Methods in Natural Language Processing, 10204–10214. Association for Computational Linguistics.
- Zouhar et al. (2023) Vilém Zouhar, Věra Kloudová, Martin Popel, and Ondřej Bojar. 2023. Evaluating optimal reference translations.